OpenStudy (anonymous):

differentiate x(2x+3)^1/2

OpenStudy (anonymous):

\[ x \sqrt{2x+3} \]Use the product rule: \[\frac{d(\sqrt{2x+3)}}{dx}x + \frac{d(x)}{dx}\sqrt{2x+3} \]

OpenStudy (anonymous):

Note that, in order to differentiate sqrt(2x + 3), we will need to use the chain rule.